OpenStudy (anonymous):

y=8/3x+1 y=-8/3x Parallel Perpendicular or Neither

OpenStudy (anonymous):

parallel line has same slope to they are paralled

OpenStudy (anonymous):

y=mx+b m is slope

OpenStudy (anonymous):

oops didn't see negative they are neither

OpenStudy (anonymous):

perpendicual lines have opposite reciprocal slopes such as 8/3 and -3/8

OpenStudy (anonymous):

slope of line one= 8/3 slope of line two=-8/3 m1 * m2 is not equal to -1 and they are not equal either. Hence they are neither parallel nor prependicular